Find the p -value for the hypothesis test. A random sample of size 50 is taken. The sample has a mean of 420 and a standard deviation of 81. H 0 : µ = 400 H a : µ > 400 The p -value for the hypothesis test is . Your answer should be rounded to 4 decimal places.
Answer #1

Test Statistic :-


t = 1.7459

Decision based on P value
P - value = P ( t > 1.7459 ) = 0.0435

Looking for t = 1.7459 in t table across n - 1 = 50 - 1 = 49 degree of freedom to find the P value for 1 tail.